# Anti-rotors and algebra invariants

### v37

@@ 2. The Skewer and the anti-rotor
Our program begins with consideration of the transformations of an algebra $A$ to itself that are dependent on its bilinear product operation, as a means of studying the algebra. Each such unary operation, $f:D_{A,f}\rightarrow A$, is a function as well as vector field (e.g., $f(s)=s^{-1}$, $f(s)=s^{2}$, etc.), where $D_{A,f}$ is an appropriate domain in the algebra’s vector space of elements. In examination of a vector field it is of interest to consider the associated infinitesimal rotation tensor (the exterior derivative of the differential 1-form dual to the vector field; its Hodge dual is the curl of the vector field). Its computation requires choice of a metric (i.e., deriving from a symmetric bilinear form, and also referred to as a scalar product or metric tensor). Specification of the standard topology entails imposition of the Euclidean inner product as the metric, allowing the computation to proceed. However, this choice of metric is arbitrary (i.e., not inherent in the axioms of an algebra), and thus can only be considered a default against which the results of all other choices should be compared. Hence, with respect to a specific choice of function $f$ to study, we are motivated to examine the mapping,
@@ 4.2. Special unital norms
(4.7) $\log\ell_{\rm sp}(s^{-1})$ $=$ $\frac{1}{\|{\bf 1}\|^{2}}\int_{\bf 1}^{s^{-1}}\left(L_{\shortparallel}t^{-1}\right)\cdot\mathbf{d}t$
$=$ $\frac{1}{\|{\bf 1}\|^{2}}\int_{\bf 1}^{s}\left(L_{\shortparallel}y\right)\cdot\mathbf{d}\left(y^{-1}\right)$
@@ 8.1. Type I anti-rotors
For $j\in\mathbb{Z}$, $j\neq 0,1$, if $s^{j}$ substitutes for $s^{-1}$ in Definition 3.1, the set of metrics satisfying the accordingly modified (3.1) is the Type I anti-rotor associated with $f(s)=s^{j}$.
@@
One might then wonder if the association of an algebra with the above multiplicity of anti-rotors (including those based on linear combinations of the above $s^{j}$) could thereby be an endless source of algebra invariants. If so, it would be appropriate to wonder how one could bring order to such a collection of processes that literally spew out invariants.
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. For each algebra, all of its Type I anti-rotors associated with the various $f(s)=s^{j}$ are identical for “small” $|j|$. Extrapolating, for any given algebra, if the above holds for all $j\in\mathbb{Z}$ except for $j=0$ or $j=1$ (by definition), linearity of the exterior derivative means it would hold for arbitrary linear combinations of the various $s^{j}$ - implying that with respect to suitable domains in $\mathbb{R}^{n}$ it would hold for functions expressible by infinite series in these powers of $s$. That is, the kernel of $\not{\text{S}}_{f}$ would be independent of $f$.
However, the behavior cited in the second sentence of the prior paragraph was unanticipated and seems surprising. There are several obvious avenues one might initially think would lead to an understanding of it.
First, let $\hat{s}$ be the left regular representation of an algebra with elements $s$, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. The behavior noted in the penultimate paragraph suggests that since $Ls$ is irrotational, and assuming $(L\hat{s}L^{-1})Ls=Ls^{2}$ is irrotational, then repeated applications of $L\hat{s}L^{-1}$ will continue to produce irrotational vector fields, i.e., $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls=Ls^{j}$ is irrotational. Why? One might suspect that this follows because, in general, when an operator applied to an irrotational vector field results in an irrotational vector field, then it will continue to produce irrotational vector fields on all successive applications - which is false. That is, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. This does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational.
Another possibility for understanding the behavior cited in the second paragraph following Corollary 8.2, is inspired by the easily appreciated truth of the generalization of that behavior to all $j\in\mathbb{Z}$ in the particular case of the full matrix algebra $M_{n}(\mathbb{R})$. Specifically, there is the well known Matrix Calculus expression valid for all $j\in\mathbb{Z}$ and for $s$ in a neighborhood of $\mathbf{1}$,
@@
where Tr indicates trace (in the Matrix Calculus literature, $s$ in (8.1) is typically written as $\mathbf{X}$). Using Jacobi’s formula, we have noted earlier that the transpose operation T (i.e., $\mathcal{T}$) is an uncurling metric for $M_{n}(\mathbb{R})$ in the exceptional $f(s)=s^{-1}$ case for which (8.1) gives no useful information. Thus, for all other $j\in\mathbb{Z}$, (8.1) indicates that T is indeed also a member (uncurling metric) of each of the Type I anti-rotors defined by the different $f(s)=s^{j}$ (since the expression on the left-hand-side of (8.1) is a gradient). So, according to an easy extrapolation of the argument in the proof of Proposition 4.6 (extending the argument from $f(s)=s^{-1}$ to $f(s)=s^{j}$, $j\neq 0,1$, to again indicate that the only uncurling metrics for $M_{n}(\mathbb{R})$ are multiples of $\mathcal{T}$), here we have an example of an algebra whose space of uncurling metrics is indeed independent of the choice of $f(s)=s^{j}$, for $j\neq 0,1$. On the other hand, it is not obvious how this observation can be adapted to the case of subalgebras of $M_{n}(\mathbb{R})$.
As an aside, the observations in the last paragraph lead to,
Proposition 8.3.
For any given algebra, the Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$, $j\neq 0,1$, has nontrivial members.
Proof.
According to the penultimate paragraph, the algebra $M_{n}(\mathbb{R})$ is such that its Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$ has nontrivial members - resulting from scalar multiples of the transpose operator. Identical arguments as appearing in the proofs of Lemma 4.7 and Theorem 4.8, except replacing $s^{-1}$ with $s^{j}$, then establish the proposition.∎
@@
Returning to our original quest, a third way to try to find an explanation for the behavior cited in the second paragraph following Corollary 8.2, is to exploit the fact that pullbacks commute with the exterior derivative. Thus, suppose it could be shown that for any $\neq 0,1, jj\in\mathbb{Z}$, one can find a $j$-dependent diffeomorphism $\phi$ between relevant domains in $\mathbb{R}^{n}$ resulting in a $j$-dependent pullback $\phi^{*}$ between the 1-forms $(\mu s^{-1})\cdot\mathbf{d}s$ and $(\mu s^{j})\cdot\mathbf{d}s$. Then we would have $\phi^{*}\big[(\mu s^{-1})\cdot\mathbf{d}s\big]=(\mu s^{j})\cdot\mathbf{d}s$, along with a similar equation expressing $\big(\mu s^{-1}\big)\cdot\mathbf{d}s$ as a pullback of $\big(\mu s^{j}\big)\cdot\mathbf{d}s$. Commutativity of the exterior derivative and a pullback would then imply that the exterior derivative of $(\mu s^{j})\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $(\mu s^{-1})\cdot\mathbf{d}s$ vanishes - and we would have our explanation.
@@
However, proving that all those pullbacks exist would involve a demonstration of the existence of solutions to each set of nonlinear partial differential equations arising from the choice of any given algebra. For individual cases of a particular algebra and a particular value of $j$, along with the stipulation that the relevant diffeomorphisms must have only rational function components, the problem might be addressed with known techniques of computational algebraic geometry. But it is difficult to overcome the impression that the general problem is intractable, i.e., that of establishing the existence of the requisite pullbacks for any algebra and any $j\neq 0,1$.
@@
To summarize: On its face, it would seem very unlikely that Conjecture 8.4 is true. And yet, one can generate a great deal of computational evidence in its favor - beginning with the algebras listed in Table III. But either way, resolution of the conjecture will lead to an interesting result.
@@ 8.2. Type II anti-rotors
We have defined the Type I anti-rotors based on different choices of powers of $s$ replacing $s^{-1}$ in Definition 3.1. We now introduce “Type II anti-rotors”, which will be generally referred to individually as “$k$th-order anti-rotors”. Each of these is a set of metrics that uncurls infinitely many vector fields given by certain powers of $s$. They are defined in a completely different manner than Type I anti-rotors.
@@
where $\hat{s}_{L}$ is the left regular representation, and $\hat{s}_{R}$ is the matrix effecting right multiplication in the given algebra. Suppose we can find nontrivial $\mu^{\prime}\in\text{Sym}_{n}$ (i.e., $\mu^{\prime}$ not the matrix with all entries being 0) such that,
@@
where in obtaining the first equality above from (8.2) we have used the characteristic property of an adjoint in the context of an inner product. Note that the exterior derivative of the 1-form operated on by $\phi^{*}$ on the left-hand-side of (8.4) is zero. Since pullbacks commute with the exterior deriviative, (8.4) implies that $\mu^{\prime}s^{3}$ is irrotational, since $\mu^{\prime}s$ is irrotational (note that for $\mu\in\text{Sym}_{n}$, $\mu s$ is always irrotational, but in general $\mu s^{3}$ will not be irrotational, since (8.3) is required to effect (8.4)). In fact, we can proceed indefinitely in this manner (i.e., next replacing $\mu^{\prime}s$ on the left-hand-side of (8.4) with $\mu^{\prime}s^{3}$, etc.) to show that (8.3) implies that $\mu^{\prime}s^{j}$ is irrotational, where $j$ is any member of the sequence defined by $a_{1}=3$ and $a_{q+1}=2a_{q}+1$, i.e., the sequence $(3,7,15,31,\dots,2^{q+1}-1,\dots)$. So, all of the sudden, we have found an infinite sequence of powers of $s$ (with exponentially increasing spacing between consecutive powers) that are all uncurled by the same metric $\mu^{\prime}$ based on the latter’s satisfaction of (8.3). In fact, again representing $\text{Sym}_{n}$ by the symmetric parametrized matrix $M$ as defined in Section 2, we can compute the collection of metrics satisfying (8.3) and term it the “Type II second-order anti-rotor”, which is a parametrized matrix that uncurls the all vector fields in the sequence $\{f_{a_{q}}(s)=s^{a_{q}}:q\in\mathbb{N}\}$ (“second order”, because the exponent in the utilized diffeomorphism $\phi(s)=s^{2}$ is 2). We could specify that the first-order anti-rotor is associated with the trivial diffeomorphism $\phi(s)=s$, thereby being the set of metrics for which $\mu^{\prime}s$ has zero curl. Thus, the first-order Type II anti-rotor would be simply $\text{Sym}_{n}$. However, we will not recognize it as a Type II anti-rotor. That is, our Type II anti-rotors begin with the second order.
@@
(8.5) $\phi^{*}[(\mu^{\prime}s)\cdot\mathbf{d}s]=(\mu^{\prime}s^{k})\cdot\mathbf{d}s^{k}=(\mu^{\prime}s^{k})\cdot\left(\sum_{q=0}^{k-1}s^{q}\,\mathbf{d}s\,s^{k-q-1}\right)\\ =(\mu^{\prime}s^{k})\cdot\left(\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)[\mathbf{d}s]\right)=\left(\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)^{T}\mu^{\prime}[s^{k}]\right)\cdot\mathbf{d}s\\ =\left((\mu^{\prime}s^{k})\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)[s^{k}]\right)\cdot\mathbf{d}s=k(\mu^{\prime}s^{2k-1})\cdot\mathbf{d}s,$
where the fifth equality requires that $\mu^{\prime}$ satisfies a more complicated equation analogous to (8.3), i.e., in place of the term $(s\,\mathbf{d}s+(\mathbf{d}s)s)$ that appears in (8.2). Specifically, the $k$th-order anti-rotor is the set of $\mu^{\prime}$ satisfying,
@@
As with the $k=2$ case in the prior paragraph, we can insert $s^{2k-1}$ on the right-hand-side of (8.5), etc., to obtain the sequence of vector fields $s^{j}$ that are each uncurled by any of the members of this $k$th-order anti-rotor. In this way we obtain infinitely many different infinite sequences of vector fields, every member of the sequence being uncurled by the same $k$th-order anti-rotor. Analogous to the second-order case above, the members of the $k$th-order anti-rotor will uncurl those vector fields $s^{j}$ with $j$ a member of the sequence defined by $a_{1}=2k-1$, $a_{q+1}=k\,a_{q}+k-1$.
Based on (8.6), we formalize this as,
@@
Let $\mathcal{D}_{k}$ be the operator which when applied to $\mathbf{d}s$ results in $\mathbf{d}s^{k}$, i.e., given by the sum on the right-hand-side of (8.6). For $k>1$, the $k$th-order Type II anti-rotor is the space of metrics each of whose members results in a symmetric operator when it left-multiplies $\mathcal{D}_{k}$.
Proposition 8.6.
@@
Proof.
The Type I anti-rotor associated with a particular $f(s)=s^{j}$ encompasses all metrics uncurling the latter vector field.∎
In view of Corollary 8.2, the above proposition also indicates that Type II anti-rotors associated with isomorphic algebras are the same up to a change of basis.
To put things in perspective, on the one hand Proposition 8.3 indicates that for a given algebra, its Type I anti-rotor associated with any $f(s)=s^{j}$, $j\neq 0,1$, has nontrivial members - although the relationships between all these Type I anti-rotors are as yet undetermined (indeed, based on the computational evidence cited above, our problem is to decide whether or not they are all identical). On the other hand, we have the $k$th-order Type II anti-rotors defined by pullbacks, which for each $k$ lead to an infinite collection of vector fields whose members are powers of $s$ that are uncurled by the same parametrized matrix of uncurling metrics. However, we have not shown that Type II anti-rotors necessarily have any nontrivial members, we have no general understanding of the connections between the Type II anti-rotors of different orders, and so far we know little more than that they are subsets of Type I anti-rotors.
But there is more to consider.
For any given odd number $j>1$, Proposition 8.6 indicates that there is a $k$th-order Type II anti-rotor that addresses the vector field $s^{j}$. But the proposition says nothing about the case where $j$ is an even number. Therefore, in the context of the diffeomorphism $\phi(s)=s^{2}$ between appropriate domains, let us examine $\phi^{*}\left[(\mu s^{2})\cdot\mathbf{d}s\right]=(\mu s^{4})\cdot\mathbf{d}(s^{2})$. If $\mu^{\prime}$ is a member of the second-order Type II anti-rotor, we have

### v38

@@ 2. The Skewer and the anti-rotor
Our program begins with consideration of the transformations of an algebra $A$ to itself that are dependent on its bilinear product operation, as a means of studying the algebra. Each such unary operation, $f:D_{A,f}\rightarrow A$, is a function as well as a vector field (e.g., $f(s)=s^{-1}$, $f(s)=s^{2}$, etc.), where $D_{A,f}$ is an appropriate domain in the algebra’s vector space of elements. In examination of a vector field it is of interest to consider the associated infinitesimal rotation tensor (the exterior derivative of the differential 1-form dual to the vector field; its Hodge dual is the curl of the vector field). Its computation requires choice of a metric (i.e., deriving from a symmetric bilinear form, and also referred to as a scalar product or metric tensor). Specification of the standard topology entails imposition of the Euclidean inner product as the metric, allowing the computation to proceed. However, this choice of metric is arbitrary (i.e., not inherent in the axioms of an algebra), and thus can only be considered a default against which the results of all other choices should be compared. Hence, with respect to a specific choice of function $f$ to study, we are motivated to examine the mapping,
@@ 4.2. Special unital norms
(4.7) $\log\ell_{\rm sp}(s^{-1})$ $=$ $\frac{1}{\|{\bf 1}\|^{2}}\int_{\bf 1}^{s^{-1}}\left(L_{\shortparallel}t^{-1}\right)\cdot\mathbf{d}t=\frac{1}{\|{\bf 1}\|^{2}}\int_{\bf 1}^{s}\left(L_{\shortparallel}y\right)\cdot\mathbf{d}\left(y^{-1}\right)$
@@ 8.1. Type I anti-rotors
For $j\in\mathbb{Z}$, $j\neq 0,1$, if $s^{j}$ substitutes for $s^{-1}$ in Definition 3.1, the set of metrics satisfying the accordingly modified (3.1) is the Type I anti-rotor associated with $f(s)=s^{j}$. Its members are called uncurling metrics.
@@
The following is established using arguments that are employed in the proofs of Lemma 4.7 and Theorem 4.8, except that $s^{-1}$ is replaced with $s^{j}$.
Proposition 8.3.
For any given algebra, the Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$, $j\neq 0,1$, is at least one-dimensional.
Proof.
Consider the well known Matrix Calculus expression valid for all $j\in\mathbb{Z}$ and for $s$ in a neighborhood of $\mathbf{1}$,
@@
where Tr indicates trace (in the Matrix Calculus literature, $s$ in (8.1) is typically written as $\mathbf{X}$). Using Jacobi’s formula, we have noted earlier that the transpose operation T (i.e., $\mathcal{T}$) is an uncurling metric for $M_{n}(\mathbb{R})$ in the exceptional $f(s)=s^{-1}$ case for which (8.1) gives no useful information. Thus, since it addresses all other $j\in\mathbb{Z}$, (8.1) then indicates that T is indeed also a member (uncurling metric) of each of the Type I anti-rotors defined by the different $f(s)=s^{j}$ (since the expression on the left-hand-side of (8.1) is a gradient). Hence, the algebra $M_{n}(\mathbb{R})$ is such that its Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$, $j\neq 0,1$, has nontrivial members - resulting from scalar multiples of the transpose operator.
Now let $\bar{A}$ be an $n$-dimensional algebra and let $A$ be its left regular representation. Given $\bar{s}\in\bar{A}$ we denote its image under the left regular representation as $s$. So, $\bar{A}$ is isomorphic to $A$, and $A$ is a subalgebra of $M_{n}(\mathbb{R})$. $\mathbf{1}_{M_{n}(\mathbb{R})}$ is the $(n\times n)$-identity matrix, and observe that as an element of $M_{n}(\mathbb{R})$, $\mathbf{1}_{A}$ is likewise the $(n\times n)$-identity matrix. However, $M_{n}(\mathbb{R})$ is $n^{2}$-dimensional, while $A$ is $n$-dimensional.
As noted above, for the algebra $M_{n}(\mathbb{R})$, $\mathcal{T}$ is an uncurling metric in the Type I anti-rotor associated with $f(s)=s^{j}$. Being an uncurling metric means that integration of $\mathcal{\mathcal{T}}s^{j}$ is path-independent in an open ball devoid of non-units centered at the multiplicative identity of $M_{n}(\mathbb{R})$. Hence, $\int_{\mathbf{1}_{M_{n}({\mathbb{R})}}}^{s}(\mathcal{T}t^{j})\cdot\mathbf{d}t$ is path-independent in the latter neighborhood and, in particular, is path-independent for paths in that neighborhood restricted to be in $A$. Suppose $\mathcal{\mathcal{T}}$ maps $A$ to $A$. Then $\int_{\mathbf{1}_{M_{n}({\mathbb{R})}}}^{s}(\mathcal{T}t^{j})\cdot\mathbf{d}t$ is not only path-independent when the domain of integration is restricted to $A$, but on such a path the integrand also remains in $A$. This means that $d((\mathcal{\mathcal{T}}s^{j})\cdot\mathbf{d}s)=0$ in some neighborhood of $\mathbf{1}_{A}$ in $A$. This implies the existence of a nontrivial uncurling metric for $A$ (i.e., an uncurling metric other than the matrix whose entries are all 0), along with all its scalar multiples.
Now suppose $\mathcal{\mathcal{T}}$ does not map $A$ to $A$. In that case, as indicated in the proof of Theorem 4.8, we can find an epimorphism $\Omega:\bar{A}\rightarrow\underline{A}$ where $\underline{A}$ has dimension $m\geq 1$ and is such that the left regular representation of $\underline{A}$ is self-adjoint as a subalgebra of $M_{m}(\mathbb{R})$. Thus, from the prior paragraph, $\underline{A}$ has a nontrivial uncurling metric. Using the same reasoning employed in Theorem 4.8, we can now use $\Omega$ to show that $\bar{A}$ itself has a nontrivial uncurling metric. All of its scalar multiples are also uncurling metrics which, along with the results of the prior paragraph and the fact that $\bar{A}$ is isomorphic to $A$, implies the proposition. ∎
So at this point, one might wonder if the association of an algebra with the above multiplicity of Type I anti-rotors (each implying a set of invariants) could thereby be an endless source of algebra isomorphism invariants. If so, it would be appropriate to wonder how one could bring order to such a collection of processes that literally spew out invariants.
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. For each algebra, simple calculations show that all of its Type I anti-rotors associated with the various $f(s)=s^{j}$ are identical for “small” $|j|$. Extrapolating, for any given algebra, if the above holds for all $j\in\mathbb{Z}$ except for $j=0$ or $j=1$ (by definition), linearity of the exterior derivative means it would hold for arbitrary linear combinations of the various $s^{j}$ - implying that with respect to suitable domains in $\mathbb{R}^{n}$ it would hold for functions expressible by infinite series in these powers of $s$. That is, the kernel of $\not{\text{S}}_{f}$ would be independent of $f$.
However, the behavior cited in the second sentence of the prior paragraph was unanticipated and seems surprising. There are a couple of obvious avenues one might initially think would lead to an understanding of it.
First, for an algebra with elements $s$, let $\hat{s}$ be the left regular representation, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. The behavior noted in the penultimate paragraph suggests that since $Ls$ is irrotational, and assuming $(L\hat{s}L^{-1})Ls=Ls^{2}$ is irrotational, then repeated applications of $L\hat{s}L^{-1}$ will continue to produce irrotational vector fields, i.e., $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls=Ls^{j}$ is irrotational. But how is that possible? One might suspect that this happens because, in general, when an operator applied to an irrotational vector field results in an irrotational vector field, then it will continue to produce irrotational vector fields on all successive applications - which is false. That is, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. This does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational.
@@
Another possibility for understanding the behavior cited in the second paragraph following the proof of Proposition 8.3 is to exploit the fact that pullbacks commute with the exterior derivative. Thus, suppose it could be shown that for any $\neq 0,1, jj\in\mathbb{Z}$, one can find a $j$-dependent diffeomorphism $\phi$ between relevant domains in $\mathbb{R}^{n}$ resulting in a $j$-dependent pullback $\phi^{*}$ between the 1-forms $(\mu s^{-1})\cdot\mathbf{d}s$ and $(\mu s^{j})\cdot\mathbf{d}s$. Then we would have $\phi^{*}\big[(\mu s^{-1})\cdot\mathbf{d}s\big]=(\mu s^{j})\cdot\mathbf{d}s$, along with a similar equation expressing $\big(\mu s^{-1}\big)\cdot\mathbf{d}s$ as a pullback of $\big(\mu s^{j}\big)\cdot\mathbf{d}s$. Commutativity of the exterior derivative and a pullback would then imply that the exterior derivative of $(\mu s^{j})\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $(\mu s^{-1})\cdot\mathbf{d}s$ vanishes - and we would have our explanation.
@@
However, even given some stipulated set of $\mu$ (e.g., as the Type I uncurling metrics associated with $f(s)=s^{-1}$) proving that all those pullbacks exist would involve a demonstration of the existence of solutions to each of the nonlinear partial differential equations arising from the choice of any given algebra and any given $j$. For individual cases of a particular algebra and a particular value of $j$, along with the stipulation that the relevant diffeomorphisms must have only rational function components, the problem might be addressed with known techniques of computational algebraic geometry. But it is difficult to overcome the impression that this kind of direct attack on the general problem of establishing the existence of the requisite pullbacks for any algebra and any $j\neq 0,1$, is unworkable.
@@
To summarize: On its face, it would seem unlikely that Conjecture 8.4 is true. And yet, one can generate a great deal of computational evidence in its favor - beginning with the algebras listed in Table III. But either way, resolution of the conjecture will lead to an interesting result.
@@ 8.2. Type II anti-rotors
We have defined the Type I anti-rotors based on different choices of powers of $s$ replacing $s^{-1}$ in Definition 3.1. We now introduce “Type II anti-rotors”, which arise individually as “$k$th-order anti-rotors”. Each $k$th-order Type II anti-rotor is a set of metrics, each of whose members uncurls each member of a set of infinitely many vector fields given by certain powers of $s$. They are defined in a completely different manner than Type I anti-rotors.
@@
where $\hat{s}_{L}$ is the left regular representation, and $\hat{s}_{R}$ is the matrix effecting right multiplication in the given algebra. Suppose we can find nontrivial $\mu^{\prime}\in\text{Sym}_{n}$ (i.e., $\mu^{\prime}$ is not the matrix with all entries being 0) such that,
@@
where in obtaining the first equality above from (8.2) we have used the characteristic property of an adjoint in the context of an inner product. Note that the exterior derivative of the 1-form operated on by $\phi^{*}$ on the left-hand-side of (8.4) is zero. Since pullbacks commute with the exterior deriviative, (8.4) implies that $\mu^{\prime}s^{3}$ is irrotational, since $\mu^{\prime}s$ is irrotational (note that for $\mu\in\text{Sym}_{n}$, $\mu s$ is always irrotational, but in general $\mu s^{3}$ will not be irrotational, since (8.3) is required to effect (8.4)). In fact, we can proceed sequentially in this manner indefinitely (i.e., next replacing $\mu^{\prime}s$ on the left-hand-side of (8.4) with $\mu^{\prime}s^{3}$, etc.) to show that (8.3) implies that $\mu^{\prime}s^{j}$ is irrotational, where $j$ is any member of the sequence defined by $a_{1}=3$ and $a_{q+1}=2a_{q}+1$, i.e., the sequence $(3,7,15,31,\dots,2^{q+1}-1,\dots)$. So, all of the sudden, we have found an infinite sequence of powers of $s$ (with exponentially increasing spacing between consecutive powers) that are all uncurled by the same metric $\mu^{\prime}$ based on the latter’s satisfaction of (8.3). In fact, again representing $\text{Sym}_{n}$ by the symmetric parametrized matrix $M$ as defined in Section 2, we can compute the collection of metrics satisfying (8.3) and term it the “Type II second-order anti-rotor”, which is a parametrized matrix that uncurls the all vector fields in the sequence $\{f_{a_{q}}(s)=s^{a_{q}}:q\in\mathbb{N}\}$ (“second order”, because the exponent in the utilized diffeomorphism $\phi(s)=s^{2}$ is 2). We could specify that the first-order anti-rotor is associated with the trivial diffeomorphism $\phi(s)=s$, thereby being the set of metrics for which $\mu^{\prime}s$ has zero curl. Thus, the first-order Type II anti-rotor would be simply $\text{Sym}_{n}$. However, we will not recognize it as a Type II anti-rotor since it is not influenced by the algebra’s bilinear product. Thus, our Type II anti-rotors begin with the second order.
@@
(8.5) $\phi^{*}[(\mu^{\prime}s)\cdot\mathbf{d}s]=(\mu^{\prime}s^{k})\cdot\mathbf{d}s^{k}=(\mu^{\prime}s^{k})\cdot\left(\sum_{q=0}^{k-1}s^{q}\,(\mathbf{d}s)\,s^{k-q-1}\right)\\ =(\mu^{\prime}s^{k})\cdot\left(\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)[\mathbf{d}s]\right)=\left(\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)^{T}\mu^{\prime}[s^{k}]\right)\cdot\mathbf{d}s\\ =\left(\mu^{\prime}\left(\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}\right)[s^{k}]\right)\cdot\mathbf{d}s=k(\mu^{\prime}s^{2k-1})\cdot\mathbf{d}s,$
where the fifth equality requires that $\mu^{\prime}$ satisfies a more complicated equation analogous to (8.3), i.e., in place of the term $(s\,\mathbf{d}s+(\mathbf{d}s)s)$ that appears in (8.2). Specifically, the $k$th-order Type II anti-rotor is the set of $\mu^{\prime}$ satisfying,
@@
As with the $k=2$ case in the prior paragraph, we can replace $\mu^{\prime}s$ on the left-hand-side of (8.5) with $\mu^{\prime}s^{2k-1}$ as appears on the right-hand-side of (8.5), etc., to obtain the sequence of vector fields $s^{j}$ that are each uncurled by any of the members of this $k$th-order anti-rotor. In this way we obtain infinitely many different infinite sequences of vector fields, every member of the sequence being uncurled by the same $k$th-order anti-rotor. Analogous to the second-order case above, the members of the $k$th-order anti-rotor will uncurl those vector fields $s^{j}$ with $j$ a member of the sequence defined by $a_{1}=2k-1$, $a_{q+1}=k\,a_{q}+k-1$.
Note that a Type I anti-rotor associated with $f(s)=s^{j}$ includes every uncurling metric for $s^{j}$. Thus, a $k$th-order Type II anti-rotor is a subset of every Type I anti-rotor associated with $f(s)=s^{j}$ for every $j$ in the latter sequence.
This all boils down to the following.
@@
Let $\mathcal{D}_{k}$ be the operator which when applied to $\mathbf{d}s$ results in $\mathbf{d}s^{k}$, i.e., $\mathcal{D}_{k}\equiv\sum_{q=0}^{k-1}\hat{s}_{L}^{q}\hat{s}_{R}^{k-q-1}$. For $k>1$, the $k$th-order Type II anti-rotor is the space of metrics each of whose members results in a symmetric operator when it left-multiplies $\mathcal{D}_{k}$.
Theorem 8.6.
@@
In view of Corollary 8.2, the above theorem also indicates that for a given $k$, the $k$th-order Type II anti-rotors associated with isomorphic algebras are the same up to a change of basis.
The significance of Theorem 8.6 is that it shows that a $k$th-order Type II anti-rotor with nontrivial members implies fundamental connections between the Type I anti-rotors arising from an infinitude of different vector fields. For example, prior to introducing Type II anti-rotors, we could say nothing about the Type I anti-rotors for any particular algebra in Table III other than they were the same for the vector fields $s^{j}$ where $|j|$ is “small” and $j$ is not 0 or 1. But for any given one of those algebras, computation of its second-order anti-rotor (which was found to be identical to the limited number of computed Type I anti-rotors) indicates that infinitely many of its Type I anti-rotors at least share subspaces of positive dimension. This represents a definite step forward.
Nevertheless, to put the present challenges in perspective, on the one hand Proposition 8.3 indicates that for a given algebra, its Type I anti-rotor associated with any $f(s)=s^{j}$, $j\neq 0,1$, has nontrivial members - although the precise relationships between all these Type I anti-rotors are as yet undetermined (indeed, based on the computational evidence cited above, our problem is to decide whether or not they are all identical). On the other hand, we have the $k$th-order Type II anti-rotors defined by pullbacks, which for each $k$ lead to an infinite collection of vector fields whose members are powers of $s$ that are uncurled by the same parametrized matrix of uncurling metrics. However, although it is easy to identify nontrivial Type II anti-rotors for specific examples (and thereby identify nontrivial relationships between the members of some infinite class of Type I anti-rotors), we have not shown that Type II anti-rotors necessarily have nontrivial members for any given algebra, we have no general understanding of the connections between the Type II anti-rotors of different orders, and in the general case we know little more than that they are subsets of Type I anti-rotors.
There is another issue to consider that we are able to dispose of now, based on the fact that the sequences defined in Theorem 8.6 all begin with an odd number. That is, for any given odd number $j>1$, Theorem 8.6 indicates that there is a $k$th-order Type II anti-rotor that addresses the vector field $s^{j}$. However, the theorem says nothing about the case where $j$ is an even number. Therefore, in the context of the diffeomorphism $\phi(s)=s^{2}$ between appropriate domains, let us examine $\phi^{*}\left[(\mu s^{2})\cdot\mathbf{d}s\right]=(\mu s^{4})\cdot\mathbf{d}(s^{2})$. If $\mu^{\prime}$ is a member of the second-order Type II anti-rotor, we have